OpenStudy (anonymous):

Solve the proportion to determine the value of x. 5x-4/4x+7 = 13/11

OpenStudy (hockeychick23):

Ok so we need to use cross multiplication for this so we do (5x-4)*11= (4x+7)*13 55x-44=52x+91 subtract 52x from both sides 55x-52x-44= 52x-52x+91 3x-44= 91 Add 44 to both sides 3x-44+44=91+44 3x= 135 Divide both sides by 3 3x/3=135/3 x= 45